OpenStudy (anonymous):

8^(y-3)=2^y i dont understand?

OpenStudy (campbell_st):

you need both sides of the equation to have the same base... the you can equate the powers and solve for y 8 = 2^3 so the equation is \[(2^3)^{ y - 3} = 2^y\] just equate the powers and solve for y

OpenStudy (anonymous):

equate them?

OpenStudy (anonymous):

Put them into an equation: 3(y-3) = y Since the base is the same on both sides

OpenStudy (anonymous):

Sorry its just not making sense to me. I multiply both sides by 3?

OpenStudy (anonymous):

One law of exponents states that \[(x^a)^b = x^{ab}\] 8^(y-3) = (2^3)^(y-3) = 2^(3y-9)

OpenStudy (anonymous):

ok so after i get 2^3y-9=2^y what do i do from there? sorry math is my worse subject

OpenStudy (anonymous):

No problem. Now since they have the same base, 2. You only need to check the exponents. 3y-9 = y

OpenStudy (anonymous):

ohhh that makes sense. so you just drop the base! and i think the answer would be 4.5? thank you:)

OpenStudy (anonymous):

Thats correct, well done!

OpenStudy (anonymous):

i dont understand either